mhchen:

Given 2 real number a,b. Prove that if a

mhchen:

I have the solution for picking a rational number r, such that a < r < b But I got lost half-way through the proof:

mhchen:

Okay the general idea is to prove that m/n is between a and b. So it replaced the rational number with integer/integer by the definition of rational number. But how do I do it for an irrational number? I know that rational + irrational = irrational So I can probably add the smallest rational number between m/n and b Such that: a < m/n < t < b

mhchen:

 \[a < \frac{m}{n} < t < b\] where \[ t = \frac{m}{n} + irrational \]

mhchen:

What if I did this:  and followed the steps in the proof... Since I know rational + irrational = irrational.

mhchen:

Wait I'm an idiot, I already proved that there exists a rational number between a and b such that a < r < b. Then: \[a - \sqrt{2} < r - \sqrt{2} < b - \sqrt{2}\] and since subtraction is closed under real numbers and since r - sqrt{2} is irrational I have real < irrational < real LOL easy

mhchen:

Then again I have to write all the steps down in case I get this on the test